specific gravity bottle weighs 14.72 g when empty, 39.74 g when filled with ater, and 44.85 g when filled with a liquid. What is the relative density of the liquid?



1
Expert's answer
2021-07-19T09:48:07-0400
"m_w=39.74-14.72=25.02\ g\\m_l=44.85-14.72=30.13\ g\\RD=\frac{30.13}{25.02}=1.204"